# A Unified Approach to Routing and Cascading for LLMs

### v1

@@ 8 Conclusion
In this work, we introduced a novel framework for routing and cascading that enabled us to propose theoretically optimal strategies for both paradigms. Furthermore, we used this theoretical analysis to propose a new paradigm for model selection, cascade routing, which combines the benefits of both routing and cascading. We showed that cascade routing can significantly outperform its baselines, especially with good quality and cost estimates. Furthermore, we found that our new cascading strategy significantly outperforms existing approaches to cascading, showing our theoretical analysis also leads to practical gains. Our work provides a theoretical foundation for model selection strategies and opens up new avenues for future research in this area, especially in the direction of obtaining more accurate quality and cost estimates.
Reproducibility Statement
All our proofs are included in the appendix and contain the exact conditions under which each statement in the main text holds. We furthermore make our code available with clear instructions on how to reproduce our results.
@@ D.1 Routerbench
D.2 Other Benchmarks
@@
We specify the exact features used for the logistic regression model that serves as the quality estimator in Section 5.2. First, we include a one-hot encoding of the various datasets in each benchmark. Furthermore, for classification, we include the probability associated with the highest class and the entropy of the class probabilities if the model has been computed. If several models have been computed, we include both whether they agree on their prediction, and the JS-divergence between their class probabilities. For open-form reasoning, we include the perplexity, number of tokens, and several quantiles of the logits if the model has been computed, in accordance with Gupta et al. [2024]. If several models have been computed, we also include whether they agree on their prediction.
@@
Appendix E Additional Results
@@
In Table 5 we report the AUC scores for the RouterBench dataset across different models and noise levels for the five-shot evaluation. Our conclusions presented in Section 5.1 remain consistent with the results presented in Table 5. However, there is one notable inconsistency: in two of the three low-noise scenarios, our cascading strategy performs worse than the threshold-based baseline cascade. In the scenario with three models, we find its cause can be found in the more difficult optimization surface for the hyperparameters of our cascading strategy. Specifically, our cascading strategy at some point starts to lose quality as cost increases. By simply setting the hyperparameters of the cascading strategy once it starts to lose quality to the ones where it obtained its highest quality, we obtain a quality of $83.35\%$ over the $83.17\%$ of the baseline cascade.
@@
Cascade (Baseline) 83.163 36383.16336383.163\,36383.163 363 78.578 93178.57893178.578\,93178.578 931 76.885 93276.88593276.885\,93276.885 932 84.271 97784.27197784.271\,97784.271 977 76.593 37676.59337676.593\,37676.593 376 73.918 09773.91809773.918\,09773.918 097 87.143 25987.14325987.143\,25987.143 259 78.603 02778.60302778.603\,02778.603 027 74.942 67274.94267274.942\,67274.942 672
Cascade (Ours) 82.678 61482.67861482.678\,61482.678 614 78.793 71278.79371278.793\,71278.793 712 77.003 99177.00399177.003\,99177.003 991 84.264 85384.26485384.264\,85384.264 853 77.195 95877.19595877.195\,95877.195 958 74.297 09674.29709674.297\,09674.297 096 86.671 10186.67110186.671\,10186.671 101 78.674 62578.67462578.674\,62578.674 625 75.075 95275.07595275.075\,95275.075 952
Cascade Routing 83.818 19783.81819783.818\,19783.818 197 78.920 28778.92028778.920\,28778.920 287 77.106 73077.10673077.106\,73077.106 730 85.506 83885.50683885.506\,83885.506 838 78.822 59678.82259678.822\,59678.822 596 76.736 93676.73693676.736\,93676.736 936 88.783 67888.78367888.783\,67888.783 678 80.881 55780.88155780.881\,55780.881 557 78.020 64878.02064878.020\,64878.020 648
[…]

### v2

@@ 7 Conclusion
In this work, we introduced a novel framework for routing and cascading that enabled us to propose theoretically optimal strategies for both paradigms. Further, we used this theoretical analysis to propose a new paradigm for model selection, cascade routing, which combines the benefits of both routing and cascading. We showed that cascade routing can significantly outperform its baselines, especially with good quality and cost estimates. We also find that our new cascading strategy significantly outperforms existing approaches to cascading, showing our theoretical analysis also leads to practical gains. Our work not only provides a theoretical foundation for model selection strategies but also opens up new avenues for future research in this area, especially in the direction of obtaining more accurate quality and cost estimates.
Impact Statement
Our work has the potential to significantly impact the field of model selection strategies. By providing a theoretical foundation for routing and cascading, we have shown that these strategies can be significantly improved by using more accurate quality and cost estimates. Cascade routing combines the strengths of both routing and cascading and offers a more flexible and effective model selection strategy. Furthermore, by underscoring the importance of quality estimation, we have highlighted a critical area for future research in model selection strategies that could lead to further improvements in this area.
@@ D.1 Routerbench
D.2 Accurate Quality Estimation
Data Split
For the SWE-Bench benchmark, we use its verified data split and divide the dataset into training and calibration subsets, with each comprising $50\%$ of the data. For the Minerva Math and LiveCodeBench benchmark, we only include the Algebra portion of Minerva Math to ensure that both benchmarks have a comparable number of samples for evaluation. Similarly, we also perform a $50\%$ split of this dataset into training and calibration sets.
Evaluation Setting
For the SWE-Bench evaluation, we analyze the performance of 10 models submitted to the benchmark’s leaderboard. The logs for these models were obtained from the official SWE-Bench repository [^3] https://github.com/swe-bench/experiments. Specifically, we evaluated the following models:
  • 20240402_sweagent_claude3opus
  • 20241007_nfactorial
  • 20240728_sweagent_gpt4o
  • 20240620_sweagent_claude3.5sonnet
  • 20241016_epam-ai-run-gpt-4o
  • 20240824_gru
  • 20241106_navie-2-gpt4o-sonnet
  • 20240820_epam-ai-run-gpt-4o
  • 20241202_agentless-1.5_claude-3.5-sonnet-20241022
  • 20241028_agentless-1.5_gpt4o
For each model, we extract the time required to complete a task to measure cost.
For LiveCodeBench and Minerva Math, we evaluate the following models:
  • Qwen-2.5-Coder-7B-Instruct
  • Qwen-2.5-Coder-1.5B-Instruct
  • Qwen-2.5-Math-7B-Instruct
  • Qwen-2.5-Math-1.5B-Instruct
We conduct experiments using version 5 of the LiveCodeBench benchmark from its official repository. For Minerva Math, we utilize the LM Evaluation Harness (Gao et al., 2024) to ensure consistent and reliable evaluation.
Cost Estimation
For SWE-Bench, the cost is defined as the time (in seconds) that a model takes to complete a task. A linear regression model is fitted to predict this cost based on the query length and, when available, the cost of running other models.
For LiveCodeBench and Minerva Math, the cost is calculated as the total number of tokens in both the query and the answer, multiplied by the size of the model (in billions of parameters). Similar to SWE-Bench, a linear model is used to predict the cost based on query length and other models’ costs.
Quality Estimation
For ex-ante quality estimation in SWE-Bench, we train a logistic regression model that predicts quality based on the query length and a one-hot encoded variable representing the query’s source repository. Post-hoc quality estimation leverages the ground-truth quality scores computed during evaluation.
For ex-ante quality estimation in Minerva Math and LiveCodeBench, we include the query length, query source (Minerva Math or LiveCodeBench), and the difficulty level of the problem as defined by the benchmark. Post-hoc quality estimation incorporates additional information, such as whether the parsed answers from different models agree with one another.
D.3 Poor Quality Estimation
@@
We specify the exact features used for the logistic regression model that serves as the quality estimator in Section 5.2. First, we include a one-hot encoding of the various datasets in each benchmark. Furthermore, for classification, we include the probability associated with the highest class and the entropy of the class probabilities if the model has been computed. If several models have been computed, we include both whether they agree on their prediction, and the JS-divergence between their class probabilities. For open-form reasoning, we include the perplexity, number of tokens, and several quantiles of the logits if the model has been computed, in accordance with Gupta et al. (2024). If several models have been computed, we also include whether they agree on their prediction.
@@
Appendix E Confidence Intervals
To check whether the results obtained by cascade routing are significantly higher than our baselines in Tables 1, 2 and 10, we perform bootstrapping on the samples in the dataset. Specifically, we compute the confidence interval associated with the difference between the AUC scores of cascade routing and the baselines. If this difference is positive and its $2\sigma$ confidence interval does not contain zero, we can conclude that cascade routing is significantly better than the baseline. These confidence intervals are reported in Tables 4, 5 and 6.
[…]
Three Models Five Models Eleven Models
Low Med High Low Med High Low Med High
Cascade Routing (Ours) $82.37^{+0.31}_{-0.32}$ $76.57^{+0.34}_{-0.35}$ $73.23^{+0.37}_{-0.38}$ $84.33^{+0.29}_{-0.29}$ $76.32^{+0.34}_{-0.37}$ $72.75^{+0.36}_{-0.40}$ $87.24^{+0.23}_{-0.26}$ $77.58^{+0.30}_{-0.33}$ $74.41^{+0.33}_{-0.36}$
-- Routing $2.64^{+0.15}_{-0.16}$ $1.59^{+0.13}_{-0.15}$ $1.40^{+0.15}_{-0.17}$ $3.10^{+0.17}_{-0.15}$ $1.88^{+0.17}_{-0.16}$ $1.41^{+0.17}_{-0.17}$ $4.00^{+0.17}_{-0.21}$ $2.94^{+0.20}_{-0.21}$ $1.73^{+0.19}_{-0.19}$
-- Cascade (Baseline) $1.50^{+0.12}_{-0.12}$ $1.91^{+0.18}_{-0.19}$ $0.74^{+0.19}_{-0.18}$ $2.00^{+0.17}_{-0.15}$ $3.29^{+0.26}_{-0.27}$ $3.22^{+0.24}_{-0.24}$ $2.76^{+0.14}_{-0.14}$ $3.92^{+0.28}_{-0.28}$ $4.61^{+0.28}_{-0.28}$
-- Cascade (Ours) $1.28^{+0.12}_{-0.11}$ $0.39^{+0.15}_{-0.14}$ $0.54^{+0.18}_{-0.17}$ $1.27^{+0.10}_{-0.11}$ $1.14^{+0.20}_{-0.21}$ $2.57^{+0.21}_{-0.26}$ $2.77^{+0.13}_{-0.13}$ $2.46^{+0.22}_{-0.24}$ $4.14^{+0.25}_{-0.27}$
[…]
Three Models Five Models Eleven Models
Low Med High Low Med High Low Med High
Cascade Routing (Ours) $83.79^{+0.3}_{-0.33}$ $78.85^{+0.29}_{-0.34}$ $77.1^{+0.32}_{-0.35}$ $85.5^{+0.26}_{-0.26}$ $78.77^{+0.32}_{-0.33}$ $76.74^{+0.34}_{-0.36}$ $88.78^{+0.22}_{-0.22}$ $80.89^{+0.28}_{-0.29}$ $78.03^{+0.3}_{-0.31}$
-- Routing $2.3^{+0.14}_{-0.13}$ $1.64^{+0.15}_{-0.15}$ $1.1^{+0.13}_{-0.14}$ $3.08^{+0.16}_{-0.14}$ $1.94^{+0.16}_{-0.14}$ $1.21^{+0.14}_{-0.15}$ $3.43^{+0.15}_{-0.17}$ $3.13^{+0.19}_{-0.2}$ $1.6^{+0.17}_{-0.16}$
-- Cascade (Baseline) $-0.64^{+0.11}_{-0.1}$ $0.28^{+0.13}_{-0.14}$ $0.22^{+0.16}_{-0.16}$ $1.23^{+0.12}_{-0.12}$ $2.19^{+0.21}_{-0.21}$ $2.83^{+0.24}_{-0.24}$ $1.64^{+0.12}_{-0.13}$ $2.29^{+0.24}_{-0.24}$ $3.09^{+0.27}_{-0.26}$
-- Cascade (Ours) $1.02^{+0.1}_{-0.09}$ $\bf{0.09^{+0.11}_{-0.11}}$ $\bf{0.1^{+0.14}_{-0.14}}$ $1.25^{+0.1}_{-0.09}$ $1.59^{+0.17}_{-0.17}$ $2.45^{+0.21}_{-0.21}$ $2.06^{+0.1}_{-0.1}$ $2.22^{+0.21}_{-0.19}$ $2.95^{+0.23}_{-0.24}$
Table 6: AUC scores on the realistic benchmarks with $2\sigma$ confidence intervals. Bold numbers indicate that the confidence interval contains zero.
SWE-Bench Math+Code Classification Open-Form
10 Models 5 Models Qwen Llama Gemma Mistral Llama Gemma Mistral
Cascade Routing (Ours) $54.21^{+7.49}_{-7.17}$ $51.20^{+7.44}_{-7.22}$ $48.51^{+2.95}_{-2.99}$ $75.56^{+1.22}_{-1.16}$ $64.89^{+1.36}_{-1.42}$ $65.02^{+1.40}_{-1.24}$ $79.95^{+1.28}_{-1.33}$ $59.70^{+1.62}_{-1.61}$ $58.77^{+1.42}_{-1.50}$
-- Routing $13.65^{+4.50}_{-4.32}$ $11.71^{+4.39}_{-4.14}$ $1.11^{+0.81}_{-0.80}$ $0.60^{+0.32}_{-0.34}$ $\bf{0.39^{+0.50}_{-0.47}}$ $\bf{0.08^{+0.12}_{-0.10}}$ $0.56^{+0.38}_{-0.42}$ $1.26^{+0.57}_{-0.55}$ $\bf{0.02^{+0.05}_{-0.05}}$
-- Cascade (Baseline) $15.36^{+3.90}_{-3.22}$ $5.17^{+1.69}_{-1.63}$ $10.77^{+1.71}_{-1.72}$ $0.71^{+0.30}_{-0.28}$ $10.51^{+0.62}_{-0.61}$ $3.79^{+0.73}_{-0.77}$ $0.65^{+0.23}_{-0.27}$ $3.47^{+0.37}_{-0.40}$ $10.45^{+1.15}_{-1.06}$
-- Cascade (Ours) $\bf{0.83^{+1.90}_{-1.50}}$ $\bf{0.11^{+1.05}_{-1.04}}$ $1.82^{+0.58}_{-0.55}$ $\bf{0.06^{+0.15}_{-0.17}}$ $2.04^{+0.33}_{-0.31}$ $1.67^{+0.41}_{-0.42}$ $0.20^{+0.19}_{-0.19}$ $2.00^{+0.25}_{-0.25}$ $3.06^{+0.65}_{-0.63}$
Appendix F Additional Experiments
F.1 Ablation Study
We conduct an ablation study to examine the impact of various design choices in cascade routing on performance and runtime. Runtime is a critical factor because the overhead introduced by the strategy must be negligible compared to the time required for model computation. If the strategy adds significant overhead, its performance gains may be offset by the increased runtime. We also include an additional ablation that specifically targets runtime on random data in Section F.2.
Table 7: AUC scores and average runtime for variations of cascade routing on RouterBench when using all eleven models.
Low-Noise Medium-Noise High-Noise
AUC (%) Time (ms) AUC (%) Time (ms) AUC (%) Time (ms)
Cascade Routing 87.288 66387.28866387.288\,66387.288 663 15.263 21915.26321915.263\,21915.263 219 77.606 26077.60626077.606\,26077.606 260 9.525 3379.5253379.525\,3379.525 337 74.408 80474.40880474.408\,80474.408 804 13.684 28513.68428513.684\,28513.684 285
Slow 87.299 14987.29914987.299\,14987.299 149 78.883 86078.88386078.883\,86078.883 860 77.608 25877.60825877.608\,25877.608 258 87.722 17387.72217387.722\,17387.722 173 74.395 60474.39560474.395\,60474.395 604 88.757 83788.75783788.757\,83788.757 837
Greedy 85.930 46585.93046585.930\,46585.930 465 1.387 4291.3874291.387\,4291.387 429 77.166 75477.16675477.166\,75477.166 754 1.173 8921.1738921.173\,8921.173 892 74.350 34974.35034974.350\,34974.350 349 0.885 6550.8856550.885\,6550.885 655
No-Expect 85.977 67585.97767585.977\,67585.977 675 4.784 6594.7846594.784\,6594.784 659 77.111 51377.11151377.111\,51377.111 513 2.492 0102.4920102.492\,0102.492 010 74.347 16374.34716374.347\,16374.347 163 2.080 3142.0803142.080\,3142.080 314
Results
Table 7 presents the results. As expected, the Slow variation is almost an order of magnitude slower while achieving similar performance. In contrast, both Greedy and No-Expect are faster but perform worse in the low- and medium-noise scenarios by $0.5\%$ to $1.3\%$. Interestingly, there is a much smaller performance gap in the high-noise scenario. This is due to the very low variance in the quality estimates, since the linear model used for quality estimation predicts an almost constant value for each query in this scenario, making the expected value computation less important.
F.2 Runtime Analysis
We further analyze the runtime of the four variants of cascade routing presented in Section F.1. Specifically, we perform experiments with random data, scaling the number of models to 80 to evaluate the runtime of all variants. Furthermore, we include a fifth variant of cascade routing in the analysis Max-Depth, which restricts cascade routing to a maximum depth of $3$ models. Max-Depth does not reduce performance of cascade routing if the optimal depth is less than or equal to $3$ models. However, it does significantly reduce the runtime of cascade routing.
For each number of models, we generate $100$ data points, each with random quality and cost estimates associated with each model. For each point, we generate the hyperparameters $\lambda_{1},...,\lambda_{k}$ and $\gamma$ randomly. We then report the average runtime of the five variants of cascade routing in Fig. 3.
Figure 3: Runtime of cascade routing variants for different numbers of models.
The results show the varying computational complexity of the different variants of cascade routing. Slow has the highest runtime, and becomes computationally too expensive even when using less than $20$ models. In contrast, standard cascade routing has a significantly lower runtime, and is able to handle up to $40$ models within a $1$ second runtime. Its faster variant, Max-Depth, is able to handle up to $80$ models within a $1$ second runtime. Furthermore, we now also see a clear difference between No-Expect and Greedy. While Greedy remains computationally very cheap even for $80$ models, No-Expect has a significantly higher runtime, even obtaining higher runtimes than Max-Depth for $80$ models.
Thus, the conclusions from Section F.1 are further supported by the runtime analysis: Greedy is the most efficient variant of cascade routing, while Normal is the most efficient variant that does not compromise performance. Max-Depth is a good choice if the optimal depth is known to be less than or equal to $3$ models, as it significantly reduces runtime without compromising performance. Since cascades of more than $3$ models are rare, Max-Depth is a good choice in practice.
Appendix G Detailed Results
We present benchmark-specific AUC values for the experiment performed in Section 5.2 in Table 8 for classification and Table 9 for open-form reasoning.
Table 8: Classification AUC values for each benchmark separately for the experiment performed in Section 5.2.
Llama Gemma Mistral
MMLU ARC MixEval MMLU ARC MixEval MMLU ARC MixEval
Linear Interp. $53.82$ $93.15$ $82.86$ $39.40$ $82.28$ $70.97$ $39.76$ $85.39$ $73.03$
Routing $55.32$ $93.12$ $82.86$ $40.01$ $83.13$ $73.12$ $40.61$ $85.64$ $74.28$
Cascade (Baseline) $54.80$ $94.08$ $84.15$ $36.43$ $77.53$ $66.10$ $36.99$ $83.88$ $72.73$
Cascade (Ours) $55.05$ $94.16$ $84.00$ $37.68$ $79.80$ $70.57$ $37.03$ $86.27$ $74.42$
Cascade Routing (Ours) $55.40$ $93.90$ $83.91$ $39.93$ $83.74$ $73.16$ $40.56$ $86.52$ $74.64$
Table 9: Open-form AUC values for each benchmark separately for the experiment performed in Section 5.2.
Llama Gemma Mistral
MMLU GSM8k MMLU GSM8k MMLU GSM8k
Linear Interp. $65.64$ $94.43$ $36.52$ $73.86$ $41.40$ $67.84$
Routing $65.75$ $94.15$ $38.08$ $75.01$ $43.03$ $68.00$
Cascade (Baseline) $66.07$ $95.17$ $35.76$ $68.44$ $38.88$ $60.82$
Cascade (Ours) $66.25$ $94.94$ $38.16$ $71.10$ $40.76$ $64.53$
Cascade Routing (Ours) $66.60$ $94.69$ $40.43$ $75.25$ $42.93$ $68.30$
Appendix H Additional Results
@@
In Table 10 we report the AUC scores for the RouterBench dataset for different noise levels for the five-shot evaluation. Our conclusions presented in Section 5.1 remain consistent with the results presented in Table 10. However, there is one notable inconsistency: in two of the three low-noise scenarios, our cascading strategy performs worse than the threshold-based baseline cascade. In the scenario with three models, we find its cause can be found in the more difficult optimization surface for the hyperparameters of our cascading strategy. Specifically, our cascading strategy at some point starts to lose quality as cost increases. By simply setting the hyperparameters of the cascading strategy once it starts to lose quality to the ones where it obtained its highest quality, we obtain a quality of $83.35\%$ over the $83.17\%$ of the baseline cascade.
@@
Cascade (Baseline) 83.163 36383.16336383.163\,36383.163 363 $78.58$ $76.89$ 84.271 97784.27197784.271\,97784.271 977 76.593 37676.59337676.593\,37676.593 376 73.918 09773.91809773.918\,09773.918 097 87.143 25987.14325987.143\,25987.143 259 78.603 02778.60302778.603\,02778.603 027 74.942 67274.94267274.942\,67274.942 672
Cascade (Ours) 82.781 41382.78141382.781\,41382.781 413 $78.77$ $77.01$ 84.259 06784.25906784.259\,06784.259 067 77.194 32977.19432977.194\,32977.194 329 74.298 90174.29890174.298\,90174.298 901 86.722 47186.72247186.722\,47186.722 471 78.671 22578.67122578.671\,22578.671 225 75.081 53075.08153075.081\,53075.081 530
Cascade Routing (Ours) 83.798 99983.79899983.798\,99983.798 999 78.858 86278.85886278.858\,86278.858 862 77.105 49377.10549377.105\,49377.105 493 85.503 88685.50388685.503\,88685.503 886 78.781 56878.78156878.781\,56878.781 568 76.750 74776.75074776.750\,74776.750 747 88.780 02588.78002588.780\,02588.780 025 80.900 84880.90084880.900\,84880.900 848 78.035 32278.03532278.035\,32278.035 322
Table 11: AUC scores on several benchmarks for the Mistral model family. Highest numbers are bolded, underlined numbers are within the $95\%$ confidence intervals of the highest number. For confidence intervals, see App. E.
Classification Open-Form
Linear Interp. 63.389 43563.38943563.389\,43563.389 435 53.859 56753.85956753.859\,56753.859 567
Routing $64.89$ $58.71$
Cascade (Baseline) 61.202 08161.20208161.202\,08161.202 081 48.288 64648.28864648.288\,64648.288 646
Cascade (Ours) 63.311 54163.31154163.311\,54163.311 541 55.507 38055.50738055.507\,38055.507 380
Cascade Routing (Ours) 64.970 63464.97063464.970\,63464.970 634 58.728 56158.72856158.728\,56158.728 561